Jack paid $13.75 to buy 3 slices of pizza and 4 hot dogs at a concession stand. Sarah paid $9.75 to buy 2 slices of pizza and 3 hot dogs from the same concession stand. What was the price of a slice of pizza?
Answer by macston(5194)

.
P=cost of piece of pizza; H=price of hot dog
.
3P+4H=$13.75 . Multiply this by 2.
2P+3H=$9.75 . Multiply this by 3.
.
6P+8H=$27.50
6P+9H=$29.25 . Subtract.
-H=-$1.75 . Multiply each side by -1.
H=$1.75
.
A hot dog costs $1.75.
.
2P+3H=$9.75
2P+3($1.75)=$9.75
2P+$5.25=$9.75
2P=$4.50
P=$2.25
.
ANSWER: A slice of pizza costs $2.25.
.
CHECK:
.
3P+4H=$13.75
3($2.25)+4($1.75)=$13.75
$6.75+$7.00=$13.75
$13.75=$13.75
.